OpenStudy (anonymous):

Find the distance between P and Q. P(sqrt 7,0); Q(0,3)

OpenStudy (debbieg):

The distance between any two points is: \(\large d=\sqrt{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}\) Can you use that with your points to compute the distance?

OpenStudy (anonymous):

@DebbieG yes but the \[\sqrt{7}\] is confusing me

OpenStudy (anonymous):

do i change it to decimal points? @DebbieG

OpenStudy (debbieg):

Nooo.... :) don't be afraid of the square root. :) Plug it into the formula, and watch what happens! It will be very easily taken care of.

OpenStudy (debbieg):

What will you get for the difference of the x's? And then you will SQUARE that. It all works out.

OpenStudy (anonymous):

@DebbieG can you pleaseee help me solve this my answers are all wrong :(

OpenStudy (anonymous):

What's your question

OpenStudy (debbieg):

Let's start by substituting your points in to the part that goes under the square root: \(\large (x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}\) You have the points :\(\large (\sqrt{7},0),(0,3)\). so those are your \(\large (x_{1},y_{1}),(x_{2},y_{2})\). Can you plug them in? Tell me what you get inside each parenthesis: \(\large (x_{1}-x_{2})=??\) \(\large (y_{1}-y_{2})=??\)